OpenStudy (anonymous):

In a game, I roll three dice. If a six appears one time, I win $1.00; if two sixes appear, I win $3.00; and if three sixes appear, I win $10.00. What is the probability that I win $1.00? Explain your answer.

OpenStudy (kropot72):

The probability of rolling a six on one die is 1/6. The probability of not rolling a six on a die is 5/6. Therefore the probability of winning $1 is \[\frac{1}{6}\times \frac{5}{6} \times \frac{5}{6}=you\ can\ calculate\]